the copper (ii) ion reacts with phosphate ion to form an insoluable ionic compound...Lap= 1.40x10..-37....
the copper (ii)ion also forms a complex ion with ammonia...(Cu(NH3)4)+2...Kf=5.03×10..13
write the chemical formula that occurs between them. also make sure you include states of matter.
Answer #1

Reaction 1 : Cu3(PO4)2 (s) 3 Cu2+ (aq) + 2 PO43- (aq) : Ksp = 1.40 x 10-37

Reaction 2 : Cu2+ (aq) + 4 NH3 (aq) Cu(NH3)42+ (aq) : Kf = 5.03 x 1013

Multiply Reaction 2 by a factor of 3 and add to Reaction 1 to cancel out the 3 Cu2+ (aq) term

Overall reaction : Cu3(PO4)2 (s) + 12 NH3 (aq) 3 Cu(NH3)42+ (aq) + 2 PO43- (aq)

overall equilibrium constant, K = (Ksp) * (Kf)3

K = (1.40 x 10-37) * (5.03 x 1013)3

K = (1.40 x 10-37) * (1.27 x 1041)

K = 1.78 x 104